OpenStudy (anonymous):

y varies inversely as twice x. When x = 2, y = 3. Find y when x = 6.

OpenStudy (blurbendy):

y = k/(2x) x = 2 y = 3 (3) = k/(2(2)) 3 = k/4 k = 12 y = 12/(2x) Plug in x = 6: y = 12/(2(6)) = 12/12 = 1